METCRAX: Logbook Entries

METCRAX: Site East_Upper Messages, 7 Entries..

Return to Logbook Contents Page
Entry Date Title Site Author #Graphics
165 Fri 03-Nov-2006eu teardownEast_UpperOncley
132 Wed 25-Oct-2006krypton kleanEast_UpperOncley
123 Tue 24-Oct-2006logger wires brokeEast_UpperOncley
81 Sun 15-Oct-2006network restored at euEast_UpperMartin
69 Fri 13-Oct-2006eu sensor serial numbersEast_UpperMartin
61 Thu 12-Oct-2006Tried to fix the problem with the 3 m RH at euEast_UpperHorst
20 Fri 06-Oct-2006radiometers cleanedEast_UpperHoch

165: LOG, Site East_Upper, Fri 03-Nov-2006 21:14:50 MST, eu teardown
Serial numbers from today's teardown of eu:	15571	31975
PSP.out	17414
PYG.out	31981
Tsurf	3577-1
Rnet-parallel	Q94125
Rnet-horiz	Q97069
Gsoil	H993561
Tsoil	NCAR002
TP01	200242
logger	9862

0.5m	CSAT	1123	TRH	6
1.5m	CSAT	364	TRH	7
5m	CSAT	1121	TRH	8	KH2O	1525
8.5m	CSAT	1124	TRH	9
GPS	71919535

132: KH2O, Site East_Upper, Wed 25-Oct-2006 14:07:31 MST, krypton klean
krypton kleaned ~1248.  MSB of counts went from 0d -> 42!

123: Radiometers, Site East_Upper, Tue 24-Oct-2006 15:26:51 MST, logger wires broke
While turning off the logger for RF testing, the power wires broke at the 
switch.  I shoved them back in, but this wiring needs to be redone.

81: Network, Site East_Upper, Sun 15-Oct-2006 17:05:24 MST, network restored at eu
The power led on the etherant was on, but the
ethenet led and the wireless led were both dark.
Login over the serial connection showed that
the systems was still running and storing locally.

Power cycle on the etherant brought it right 
back and on the network.

This etherant went offline at around the time
that thinderstorms were in the area. We had
a similar experience a few days ago with the etherant 
at flr. Sebastion reported that they had lost
a couple of data loggers during the same period,probably 
due to lightning.
69: Station_Visit, Site East_Upper, Fri 13-Oct-2006 17:29:40 MST, eu sensor serial numbers
Q7		Q94125
Q7horiz		Q97069
pygin		EP31981
pygout		EP31975
pyrin		EP17414
pyrout		EP15571
Tsoil		NCAR002
HFT		H993561
TP01		200242

61: TRH, Site East_Upper, Thu 12-Oct-2006 17:25:39 MST, Tried to fix the problem with the 3 m RH at eu
Tried to fix the problem with the RH at 3m at east upper, but it appeared to be behaving itself while we were there...
20: Radiometers, Site East_Upper, Fri 06-Oct-2006 10:50:35 MST, radiometers cleaned
03.10. 15:21 cleaned 4-component radiometers, not net radiometers